# Making it to First: The Random Access Problem in DNA Storage

### v1

@@ 4 General Construction
    If the column $\bf g$ has weight $1$ in row $j$, then we have collected the vertex labeled $V_{j}$ in the graph $K_{k}$.
@@
For shorthand, we will refer to the multiset of collected edges and vertices from $K_{k}$ using the notation $\mathscr{T}$.
@@
For the case of $k=3$, the top left graph in Figure 1 represents the setup where the column with a single 1 in the first position is collected so that $\mathscr{T}=\{V_{1}\}$. The graph in the bottom left represents the event where three columns are collected where one column has support $\{1,2\}$, another column has support $\{2,3\}$, and a third column has support $\{1,3\}$ so that $\mathscr{T}=\{E_{1,2},E_{2,3},E_{1,3}\}$.
Figure 1: Recovery sets over $K_{3}$ for information strand $1$.
For convenience, we will refer to the scenario where a vertex is collected (i.e., where $V_{j}\in\mathscr{T}$) as a cycle of length $1$, and when an edge is collected twice as a cycle of length $2$. Using this interpretation, we can more precisely state conditions that allow for the recovery of any information strand in $G_{k}(x,y)$ or equivalently any vertex in $K_{k}$.
@@
Figure 2: Recovery sets over $K_{4}$ for information strand $1$.
For general $k$, $x$, and $y$, it seems to be hard to explicitly compute $T_{\max}(G_{k}({x,y}))$. However, in [17] the authors computed $\lim_{x\to\infty}T_{\max}(G_{3}({x,x}))$.
4.1 Three Information Strands
In [17], the authors provided a construction of a generator matrix that is recovery complete for $k=3$. They provided an analysis of the asymptotic behavior of the random access expectation of a recovery complete generator matrix, for the case where the multiplicity of the information strands was the same as the one of strands per edge (i.e., $y=x$). They showed that the random access expectation in this case is bounded from above by $0.8822\cdot 3$. However, if we allow the multiplicity to be any value proportional to $x$ (i.e., $y=\alpha x$) we can achieve a strictly smaller value for the expectation, as we show below. The analysis follows the same steps as in [17], and the details can be found in Appendix A.
Theorem 3.
It holds that $\lim_{x\longrightarrow{}\infty}T_{\max}(G_{3}({x,\alpha x}))\leq 3-\frac{\alpha}{3+3\alpha}-\frac{2+10\alpha+5\alpha^{2}}{9(1+\alpha)^{2}}+\frac{(1+2\alpha)^{2}(1+2\alpha)}{9(1+\alpha)^{2}(2+\alpha)}+\frac{2\alpha^{2}(9+7\alpha)}{9(1+\alpha)^{2}(3+2\alpha)^{2}}.$
By numerical analysis we obtain that the value of $\alpha$ for which the minimum is attained is approximately $0.834$. By plugging $\alpha=0.834$ to Theorem 3 we obtain the following corollary, which slightly improves the result from [17].
@@ 4.1 Three Information Strands
4.2 Four Information Strands
In this subsection, we analyze the value of $T_{\max}(G_{4}({x,y}))$ using Lemma 1. In order to apply Lemma 1, we first need to find the values of $\alpha_{i}^{s}(G_{4}({x,y}))$ for all $s\in[6x+4y-1]$ and $i\in[k]$.
@@ 4.2 Four Information Strands
    $\alpha_{i}^{3}(G_{4}({x,y}))=3\left({3x+2y\choose 3}-{x+2y\choose 3}\right)-3({x\choose 3}+{x\choose 2}y+x{y\choose 2})+3x{x+2y\choose 2}+y{6x+3y\choose 2}+{y\choose 2}(6x+3y)+{y\choose 3};$
@@
    $\alpha_{i}^{s}(G_{4}({x,y}))={6x+4\choose s}-{3x+3y\choose s}-{x+2y\choose s-1}$  for $4\leq s\leq 3x+3y$;
@@
Suppose we want to recover ${\bf e}_{1}$.
@@
In order to prove (ii), we need to count the number of 2-sets that contain ${\bf e}_{1}$ in their span. If none of the two columns is ${\bf e}_{1}$, then to recover ${\bf e}_{1}$, the 2-set must span an edge of the form $E_{1,j}$, and there are $3({x\choose 2}+x\cdot y)$ such 2-sets. Moreover, there are a total of (${y\choose 2}+y(6x+3y)$ $2$-sets that contain ${\bf e}_{1}$.
For (iii) we start by counting 3-sets that do not contain ${\bf e}_{1}$. Consider a $3$-set which lies in the span of $\langle{\bf e}_{1},{\bf e}_{j},{\bf e}_{j^{\prime}}\rangle$ for some $j,j^{\prime}\in\{2,3,4\}$. The only such $3$-sets that do not recover ${\bf e}_{1}$ are the ones where all the columns lie in $\langle{\bf e}_{j},{\bf e}_{j^{\prime}}\rangle$ and there are (${x+2y\choose 3}$ such 3-sets. Note that in this way, every such $3$-set which lies in $\langle{\bf e}_{1},{\bf e}_{j}\rangle$ was counted twice. For example if $j=2$, then the 3-sets contained in $\langle{\bf e}_{1},{\bf e}_{2}\rangle$ were counted as part of the 3-sets in $\langle{\bf e}_{1},{\bf e}_{2},{\bf e}_{3}\rangle$ but also as part of the 3-sets in $\langle{\bf e}_{1},{\bf e}_{2},{\bf e}_{4}\rangle$. There are (${x\choose 3}+{x\choose 2}y+x{y\choose 2}$ such sets (where we counted them depending on how many ${\bf e}_{2}$ they contain, which is 1, 2, or 3). The other $3$-sets that retrieve ${\bf e}_{1}$ but do not contain ${\bf e}_{1}$ are the sets where two columns lie in $\langle{\bf e}_{1},{\bf e}_{j}\rangle\setminus\{{\bf e}_{j}\}$ and the other column lies in $\langle{\bf e}_{i},{\bf e}_{i^{\prime}}\rangle$ for $i,i^{\prime}\neq j$, and there are $3x{x+2y\choose 2}$ such 3-sets. We are left to add the number of $3$-sets which contain ${\bf e}_{1}$ and there are $y{6x+3y\choose 2}+{y\choose 2}(6x+3y)+{y\choose 3}$ of them, where we counted them depending on how many ${\bf e}_{1}$ they contain, which is 1, 2, or 3.
(iv) We count the number of $s$-sets that do not recover ${\bf e}_{1}$. If the first coordinate in all the columns of the $s$-set is 00, then it does not recover ${\bf e}_{1}$, and there are (${3x+3y\choose s}$ such sets. The other case is when there is one column from $E_{1,i}$ for some $i\in\{2,3,4\}$, and the other $s-1$ columns must be contained in ${\bf e}_{j}\cup{\bf e}_{j^{\prime}}\cup E_{j,j^{\prime}}$ where $j,j^{\prime}\notin\{1,i\}$, and there are $3x{x+2y\choose s-1}$ such sets.
@@
Despite having closed expressions for the $\alpha_{i}^{s}(G_{4}({x,y}))$ for all $s\in[6x+4y-1]$ and $i\in[k]$, and thus for $T_{\max}(G_{4}({x,y}))$ (see Lemma 1), it is challenging to compute the exact value of $T_{\max}(G_{4}({x,y}))$. Nonetheless, from experiments with the computer algebra program magma, we obtain the plot shown in Figure 3, as well as the result stated in Proposition 1.
000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.811111.21.21.21.21.41.41.41.41.61.61.61.61.81.81.81.822220.850.850.850.850.870.870.870.870.890.890.890.890.910.910.910.910.930.930.930.93Ratio $\alpha=y/x$$x=5$$x=10$$x=100$$x=1000$

### v2

@@ 4.2 The Random Access Expectation of the Construction
    If the column $\bf g$ has weight $1$ with support $j$, then we have collected the vertex labeled $V_{j}$ in the graph $K_{k}$.
@@
For shorthand, we denote by $\mathscr{T}$ the multiset of collected vertices and edges from $K_{k}$.
@@
For the case of $k=3$, the top-left graph in Figure 2 illustrates the scenario where a column with a single 1 in the first coordinate is collected, corresponding to $\mathscr{T}=\{V_{1}\}$. The bottom-left graph represents the event where three columns are collected: one with support $\{1,2\}$, another with support $\{2,3\}$, and the third with support $\{1,3\}$, resulting in $\mathscr{T}=\{E_{1,2},E_{2,3},E_{1,3}\}$.
Figure 2: Recovery sets over $K_{3}$ for information strand 11.
For convenience, we will refer to the scenario where a vertex is collected (i.e., where $V_{j}\in\mathscr{T}$) as a cycle of length $1$, and when an edge is collected twice (when the two columns representing this edge are not collinear) as a cycle of length $2$. Using this interpretation, we can now state precise conditions under which any information strand in $G_{k}(x,y)$, or equivalently, any vertex of $K_{k}$, can be recovered.
@@
Proof.
Assume we want to recover the vertex $V_{i}$. By the recovery-completeness property, every vertex on the cycle is recoverable. Let $P$ be a path in $\mathscr{T}$ from $V_{i}$ to the cycle, and let $u$ be the last vertex of $P$ that lies on the cycle. Using $u$ and the edge of $P$ incident to $u$, we can recover the preceding vertex on $P$. Repeating this backtracking step recovers each earlier vertex along $P$ until we recover $V_{i}$.
Conversely, if $V_{i}$ does not belong to a connected component of $\mathscr{T}$ that contains a cycle, then it cannot be recovered. ∎
@@
Figure 3: Recovery sets over $K_{4}$ for information strand 11.
For general values of $k$, $x$, and $y$, explicitly computing $T_{\max}(G_{k}({x,y}))$ is quite challenging. However, in [19], the authors managed to derive an upper bound on $\lim_{x\to\infty}T_{\max}(G_{3}({x,x}))$, which at the time gave the best upper bound on $T(3)$. Subsequently, [6] determined the exact limit of $\lim_{x\to\infty}T_{\max}(G_{3}({x,\alpha x}))$, therefore improving the result of [19]. As [6] was published after the first draft of this paper appeared online, we include in the next subsection our own improved upper bound on $\lim_{x\to\infty}T_{\max}(G_{3}({x,\alpha x}))$ (which sharpened [19]) and, for completeness, we also summarize the results of [6].
4.2.1 Three Information Strands
In [19], the authors presented a construction of $G_{3}(x,y)$ based on a geometric approach. They analyzed the asymptotic behavior of the random access expectation for the case in which the multiplicity of the information strands equals that of the strands per edge, (i.e., $y=x$). In this setting, they established an upper bound of $0.8822\cdot 3$ on the random access expectation. However, by allowing the multiplicity to scale proportionally with $x$, i.e., $y=\alpha x$, a strictly smaller expectation can be achieved. Our analysis, which follow the same steps as in [19], shows this improvement and is proven in Appendix A.
Lemma 5.
It holds that $\lim_{x\to{}\infty}T_{\max}(G_{3}({x,\alpha x}))\leq 3-\frac{\alpha}{3+3\alpha}-\frac{2+10\alpha+5\alpha^{2}}{9(1+\alpha)^{2}}+\frac{(1+2\alpha)^{2}(1+2\alpha)}{9(1+\alpha)^{2}(2+\alpha)}+\frac{2\alpha^{2}(9+7\alpha)}{9(1+\alpha)^{2}(3+2\alpha)^{2}}.$
A numerical optimization shows that the minimum of this expression is attained at approximately $\alpha=0.834$. Substituting this value into Theorem 5 yields the following corollary, which slightly improves the result from [19].
@@ 4.2.1 Three Information Strands
In a recent paper [6], the authors managed to evaluate the exact expression of the expectation and proved that
$\lim_{x\to{}\infty}T_{\max}(G_{3}({x,\alpha x}))=\frac{153+543\alpha+805\alpha^{2}+611\alpha^{3}+234\alpha^{4}+36\alpha^{5}}{3(1+\alpha)^{2}(2+\alpha)(3+2\alpha)^{2}}.$
Their numerical optimization showed that the optimal value of $\alpha$ is $\alpha^{*}\approx 0.833968$ and for that value $\lim_{x\to{}\infty}T_{\max}(G_{3}({x,\alpha^{*}x}))\approx 0.881542\cdot 3$.
4.2.2 Four Information Strands
In this subsection, we obtain a closed form expression to the value of $T_{\max}(G_{4}({x,y}))$ using Lemma 1. To apply Lemma 1, we first compute the values of $\alpha_{i}^{s}(G_{4}({x,y}))$ for all $s\in[6x+4y-1]$ and $i\in[k]$, as established in the next lemma.
@@ 4.2.2 Four Information Strands
    $\alpha_{i}^{3}(G_{4}({x,y}))=3\left({3x+2y\choose 3}-{x+2y\choose 3}-2x\binom{y}{2}\right)-3({x\choose 3}+{x\choose 2}y+x{y\choose 2})+3x(xy+\binom{x}{2})+y{6x+3y\choose 2}+{y\choose 2}(6x+3y)+{y\choose 3}$;
@@
    $\alpha_{i}^{s}(G_{4}({x,y}))={6x+4y\choose s}-{3x+3y\choose s}-3x{x+2y\choose s-1}-3x^{2}\binom{y}{s-2}{-6x^{2}\binom{y}{s-2}}$  for $4\leq s\leq 3x+3y$;
@@
Suppose we want to recover ${\bf e}_{1}$. Recall that $E_{i,j}$ denotes all the columns that lie in $\langle{\bf e}_{i},{\bf e}_{j}\rangle\setminus\{{\bf e}_{i},{\bf e}_{j}\}$.
@@
In order to prove (ii), we count the number of $2$-sets that contain ${\bf e}_{1}$ in their span. If neither of the two columns is ${\bf e}_{1}$, then in order to recover ${\bf e}_{1}$, the $2$-set must span an edge of the form $E_{1,j}$, and there are $3({x\choose 2}+x\cdot y)$ such $2$-sets. Moreover, there are a total of ${y\choose 2}+y(6x+3y)$ 22-sets that contain ${\bf e}_{1}$.
For (iii), we start by counting the $3$-sets that do not contain ${\bf e}_{1}$. Consider a $3$-set which lies in the span of $\langle{\bf e}_{1},{\bf e}_{j},{\bf e}_{j^{\prime}}\rangle$ for some $j,j^{\prime}\in\{2,3,4\}$. The only such 3-sets that do not recover ${\bf e}_{1}$ are: a) those where all columns lie in $E_{j,j^{\prime}}$, there are $x+2y\choose 3$} such 3-sets, and b) those where one column lies in $E_{1,j}$ and there are two copies of ${\bf e}_{j^{\prime}}$, there are $2x\binom{y}{2}$ such $3$-sets. Note that in this way, every such $3$-set which lies in $\langle{\bf e}_{1},{\bf e}_{j}\rangle$ was counted twice. For example if $j=2$, then the 3-sets contained in $E_{1,2}\cup\{e_{2}\}$ were counted as part of the 3-sets in $\langle{\bf e}_{1},{\bf e}_{2},{\bf e}_{3}\rangle$ but also as part of the 3-sets in $\langle{\bf e}_{1},{\bf e}_{2},{\bf e}_{4}\rangle$. There are ${x\choose 3}+{x\choose 2}y+x{y\choose 2}$ such sets.
The remaining $3$-sets that recover ${\bf e}_{1}$ but do not contain ${\bf e}_{1}$ are the sets where one column lies in $E_{i,i^{\prime}}$ for $i,i^{\prime}\neq 1$, and the other two columns lie in $\langle{\bf e}_{1},{\bf e}_{j}\rangle\setminus\{{\bf e}_{1}\}$, for $j\neq i,i^{\prime}$, there are $3x(xy+\binom{x}{2})$ such 3-sets. Finally, we add the number of $3$-sets which contain ${\bf e}_{1}$, which is $y{6x+3y\choose 2}+{y\choose 2}(6x+3y)+{y\choose 3}$.
(iv) We count the number of $s$-sets that do not recover ${\bf e}_{1}$. If the first coordinate in all the columns of the $s$-set is 0, then it does not recover ${\bf e}_{1}$. There are $3x+3y\choose s$} such sets. The other cases are: a) when there is one column from $E_{1,i}$ for some $i\in\{2,3,4\}$, and the other s1s-1 columns must be contained in ${\bf e}_{j}\cup{\bf e}_{j^{\prime}}\cup E_{j,j^{\prime}}$ where $j,j^{\prime}\notin\{1,i\}$, there are $3x{x+2y\choose s-1}$ such sets, b) when there is one column from $E_{1,i}$, another column from $E_{1,i^{\prime}}$ ($i\neq i^{\prime}$) and s2s-2 copies of $e_{j}$ ($j\neq i,i^{\prime}$), there are $3x^{2}\binom{y}{s-2}$ such $s$-sets, and the last one c) then there is one column from $E_{1,i}$, another column from $E{i,i^{\prime}}$ and s2s-2 copies of $e_{i^{\prime\prime}}$, there are $6x^{2}\binom{y}{s-2}$ such $s$-sets.
@@
Although we have closed-form expressions for $\alpha_{i}^{s}(G_{4}({x,y}))$ for all $s\in[6x+4y-1]$ and $i\in[k]$, and therefore an explicit expression for $T_{\max}(G_{4}({x,y}))$ (see Lemma 1), computing its exact value remains challenging. Nevertheless, using the computer algebra system magma, we obtain the plot shown in Figure 4, along with the result presented in Proposition 1.
00.20.20.40.40.60.60.80.8111.21.21.41.41.61.61.81.8220.860.860.880.880.90.90.920.92Ratio $\alpha=y/x$x=5x=5x=10x=10x=100x=100x=1000x=1000
@@
Thus far, we have studied the random access expectation of our construction for $k=3$ and $k=4$, in the next subsection, we extend the analysis to general $k$, using the complete graph setup.
4.2.3 The Asymptotic Expectation for General $k$
As observed in Lemma 6, explicitly calculating the $\alpha$’s can be very challenging, especially when $k$ increases. Therefore, in this section, we derive a closed-form expression for the asymptotic random access expectation of Construction 1 (i.e., when the number of columns goes to infinity), for every $k$. To achieve this, we look at the problem from a graph viewpoint and apply Claim 3 to evaluate the random access expectation.
Assume we collect either edges or vertices from the graph $K_{k}$. Each edge is collected with probability $p=\frac{x}{ky+\binom{k}{2}x}$, and each vertex with probability $P=\frac{y}{ky+\binom{k}{2}x}$. The random access expectation of the first strand is then:
$E[\tau_{1}]=1+(1-P)+\sum_{r=2}^{\infty}\text{Pr}\left(\tau_{1}>r\right).$ (17)
The key assumption that allows us to compute the asymptotic expectation is the following observation.
Observation 1.
As we are interested in the asymptotic behavior of the random access expectation, we can assume that the same copy of an edge cannot be collected twice. This assumption is justified because, by taking $x$ to infinity, we can make the probability of collecting the same copy twice arbitrarily small.
Next, we decompose the last term in Equation (17) into disjoint cases. For that purpose, let $H_{r}$ denote the multi-set of components from $K_{k}$ collected at round $r$. Then, vertex $1$ (or information strand $1$) cannot be recovered at round $r\geq 2$ only if one of the following conditions holds:
(i) $H_{r}$ does not contain any components that involve vertex $1$, or
(ii) Vertex $1$ belongs to a tree containing $\ell+1\leq k$ vertices.
In the next lemma, we calculate the expectation of each of the cases.
Theorem 3.
Under the assumption that the same copy of an edge cannot be collected twice, we have:
${\mathbb{E}}\left(\text{ Case }i\ \right)=\frac{\left(1-(k-1)p-P\right)^{2}}{(k-1)p+P},$
${\mathbb{E}}\left(\text{ Case }ii\ \right)=(k-1)p\frac{v(2-v)}{(1-v)^{2}}+\sum_{\ell=2}^{k-1}\binom{k-1}{\ell}(\ell+1)^{\ell-1}\ell!p^{\ell}\frac{1}{\left(1-u\right)^{\ell+1}},$
where $v=\binom{k-2}{2}p+(k-2)P$ and $u=\binom{k-\ell-1}{2}p+(k-\ell-1)P$.
In particular, it holds that:
$T(k)\leq\min\limits_{\begin{subarray}{c}p,P:\\ kP+\binom{k}{2}p=1\end{subarray}}1+(1-P)+{\mathbb{E}}\left(\text{ Case }i\ \right)+{\mathbb{E}}\left(\text{ Case }ii\ \right).$ (18)
Proof.
We start with the first event:
${\mathbb{E}}\left(\text{ Case }i\ \right)$ $=\sum_{r=2}^{\infty}\left(1-(k-1)p-P\right)^{r}$
$=\frac{\left(1-(k-1)p-P\right)^{2}}{(k-1)p+P}.$
For every $r\geq 2$, this represents the probability that we don’t collect the first vertex and that no edges which are collected are adjacent to the first vertex. Since there are k1k-1 edges that are adjacent to vertex $1$ in the graph $K_{k}$, the result follows.
For the second term:
${\mathbb{E}}\left(\text{ Case }ii\ \right)=\sum_{r=2}^{\infty}\sum_{\ell=1}^{\min\{k-1,r\}}\binom{k-1}{\ell}\left(\ell+1\right)^{\ell-1}\ell!\binom{r}{\ell}p^{\ell}\left(\binom{k-\ell-1}{2}p+(k-\ell-1)P\right)^{r-\ell},$ (19)
this addresses the scenario where vertex $1$ has not been collected and where vertex $1$ belongs to a tree containing $\ell+1\leq k$ vertices, which means that the tree has at most $\ell\leq k-1$ edges. Since the number of draws is $r$, we must also have that $\ell\leq r$. In the expression, we partition the collected components in $H_{r}$ into two disjoint multi-sets:
(a) The first multi-set contains the vertex $1$ and, by assumption, forms a tree on $\ell+1$ vertices.
(b) The second multi-set contains components which are not connected to the tree in the first set.
We begin by analyzing the number of ways to select the first multiset. There are $\binom{k-1}{\ell}$ ways to choose $\ell$ additional vertices from the remaining k1k-1 (excluding vertex 1), to form a group of $\ell+1$ vertices containing vertex $1$. By Cayley’s formula, there are $\left(\ell+1\right)^{\ell-1}$ distinct trees that can be generated given $\ell+1$ vertices. Each of these trees has $\ell$ edges, which may be collected in $(\ell)!$ different orders. Since each edge is collected independently with probability $p$, this gives the contribution of the $p^{\ell}$ term.
For the second multiset, consider the $k-\ell-1$ not included in the first tree. Among these, there are $\binom{k-\ell-1}{2}$ possible edges, and each vertex can also be collected independently. Therefore, the contribution from these components over the remaining $r$-$\ell$ rounds is: $\left(\binom{k-\ell-1}{2}p+(k-\ell-1)P\right)^{r-\ell}$. Note that when $\ell=k-1$, this term vanishes, as there are no remaining vertices.
Next, we get rid of the infinite sum in equation (19). For shorthand let
$u\doteq\binom{k-\ell-1}{2}p+(k-\ell-1)P,$
$v\doteq\left(\binom{k-2}{2}p+(k-2)P\right).$
Clearly, $|u|,|v|<1$. Then, by reorganizing terms, we have
${\mathbb{E}}\left(\text{ Case }ii\ \right)$ $=\sum_{r=2}^{\infty}\sum_{\ell=1}^{\min\{k-1,r\}}\binom{k-1}{\ell}\left(\ell+1\right)^{\ell-1}\ell!\binom{r}{\ell}p^{\ell}\left(\binom{k-\ell-1}{2}p+(k-\ell-1)P\right)^{r-\ell}$
$=\sum_{r=2}^{\infty}(k-1)rp\left(\binom{k-2}{2}p+(k-2)P\right)^{r-1}+\sum_{r=2}^{\infty}\sum_{\ell=2}^{\min\{k-1,r\}}\binom{k-1}{\ell}\left(\ell+1\right)^{\ell-1}\ell!\binom{r}{\ell}p^{\ell}u^{r-\ell}$
$=(k-1)p\frac{v(2-v)}{(1-v)^{2}}+\sum_{r=2}^{\infty}\sum_{\ell=2}^{\min\{k-1,r\}}\binom{k-1}{\ell}\left(\ell+1\right)^{\ell-1}\ell!\binom{r}{\ell}p^{\ell}u^{r-\ell}$
$=(k-1)p\frac{v(2-v)}{(1-v)^{2}}+\sum_{\ell=2}^{k-1}\binom{k-1}{\ell}(\ell+1)^{\ell-1}\ell!p^{\ell}\sum_{r=\ell}^{\infty}\binom{r}{\ell}u^{r-\ell}$
$=(k-1)p\frac{v(2-v)}{(1-v)^{2}}+\sum_{\ell=2}^{k-1}\binom{k-1}{\ell}(\ell+1)^{\ell-1}\ell!p^{\ell}\frac{1}{\left(1-u\right)^{\ell+1}},$
as desired. ∎
Setting $k=4$ and $\alpha=0.95$ (the value of $\alpha=y/x$ that we used in Proposition 1) in Theorem 3 gives the upper bound $T(4)\leq 0.86375\cdot 4$. This bound is slightly stronger than the bound in Proposition 1, as expected, since Theorem 3 assumes that the same copy of an edge cannot be collected more than once. Nonetheless, as $x\to\infty$, the probability of resampling the same copy of an edge more than once tends to zero, and thus the two bounds converge.
1120204040606080801001000.790.790.80.80.810.810.820.82Ratio $\alpha=y/x$k=100k=100k=500k=500k=1000k=1000k=5000k=5000
Figure 5: Normalized (by $k$) upper bound on the random access expectation $T(k)$ for different values for $k$ and ratios $\alpha=y/x$ from Theorem 3.
Now, by substituting different values of $p$ and $P$ in Theorem 3, we can derive an upper bound on $T(k)$, and consequently on $\liminf_{k\to\infty}\frac{T(k)}{k}$. Computations in Magma suggest that the optimum of the expression in equation 18, is achieved when $p\neq P$ (as can be seen in Figure 5). Unfortunately, such asymmetric choices make the analysis cumbersome. To keep the calculations tractable, we adopt the symmetric choice $p=P=\frac{1}{k+\binom{k}{2}}=\frac{2}{k^{2}+k}$. With this choice the forthcoming corollary yields the best upper bounds currently known for both $T(k)$ and $\liminf_{k\to\infty}\frac{T(k)}{k}$.
Corollary 3.
It holds that:
$T(k)$ $\leq 1+1-\frac{2}{k^{2}+k}+\frac{(k-1)^{2}}{2(k+1)}$
$+(k-1)\frac{2}{k^{2}+k}\cdot\frac{k^{2}-3k+2}{4k-2}\cdot\frac{k^{2}+5k-2}{4k-2}$
$+\sum_{\ell=2}^{k-1}\binom{k-1}{\ell}\frac{1}{(\ell+1)^{2}}\ell!\frac{2^{\ell}k(k+1)}{(2k-\ell)^{\ell+1}}.$
Moreover:
$\liminf_{k\to\infty}\frac{T(k)}{k}\leq\frac{\pi^{2}}{12}\approx 0.822467.$
Proof.
For $p=P=\frac{1}{k+\binom{k}{2}}=\frac{2}{k^{2}+k}$ we have by Lemma 3:
$T(k)\leq 1+1-\frac{2}{k^{2}+k}+{\mathbb{E}}\left(\text{ Case }i\ \right)+{\mathbb{E}}\left(\text{ Case }ii\ \right).$
We now evaluate the expectation of both cases when $p=P=\frac{2}{k^{2}+k}$. We start with the first case.
${\mathbb{E}}\left(\text{ Case }i\ \right)=\frac{(1-kp)^{2}}{kp}=\frac{(\frac{k-1}{k+1})^{2}}{\frac{2}{k+1}}=\frac{(k-1)^{2}}{2(k+1)}.$
Which goes to $k/2$ when $k\to\infty$.
We now move on to case $ii$, we have:
$v=\binom{k-2}{2}p+(k-2)P=(k-2)p(\frac{k-3}{2}+1)=\frac{(k-2)(k-1)}{k(k+1)}.$
Thus, the first term of ${\mathbb{E}}\left(\text{ Case }ii\ \right)$ when $p=P$ equals to
$(k-1)p\frac{v(2-v)}{(1-v)^{2}}$ $=(k-1)\frac{2}{k^{2}+k}\cdot\frac{k^{2}-3k+2}{k^{2}+k}\cdot\frac{k^{2}+5k-2}{k^{2}+k}\cdot\frac{(k^{2}+k)^{2}}{(4k-2)^{2}}$
$=(k-1)\frac{2}{k^{2}+k}\cdot\frac{k^{2}-3k+2}{4k-2}\cdot\frac{k^{2}+5k-2}{4k-2},$
which goes to $\frac{1}{8}k$ when $k\to\infty$. In addition, similar to $v$, we have that
$u=\frac{(k-\ell-1)(k-\ell)}{k(k+1)},$
$1-u=\frac{2k+2\ell k-\ell^{2}-\ell}{k^{2}+k}.$
Hence, the second term of ${\mathbb{E}}\left(\text{ Case }ii\ \right)$ when $p=P$, equals to:
$\sum_{\ell=2}^{k-1}\binom{k-1}{\ell}(\ell+1)^{\ell-1}\ell!p^{\ell}\frac{1}{\left(1-u\right)^{\ell+1}}$
$=$ $\sum_{\ell=2}^{k-1}\binom{k-1}{\ell}(\ell+1)^{\ell-1}\ell!\frac{2^{\ell}}{k^{\ell}(k+1)^{\ell}}\frac{k^{\ell+1}(k+1)^{\ell+1}}{\left((\ell+1)(2k-\ell)\right)^{\ell+1}}$
$=$ $\sum_{\ell=2}^{k-1}\binom{k-1}{\ell}\frac{1}{(\ell+1)^{2}}\ell!\frac{2^{\ell}k(k+1)}{(2k-\ell)^{\ell+1}}.$
Together with what we did earlier, this gives us the required bound on $T(k)$. To obtain the bound on $\liminf_{k\to\infty}\frac{T(k)}{k}$, we analyze the asymptotic behavior of the normalized above expression when $k\to\infty$.
$\lim_{k\to\infty}\frac{1}{k}\sum_{\ell=2}^{k-1}\binom{k-1}{\ell}\frac{1}{(\ell+1)^{2}}\ell!\frac{2^{\ell}k(k+1)}{(2k-\ell)^{\ell+1}}=$
$\lim_{k\to\infty}\sum_{\ell=2}^{k-1}\frac{1}{k}\frac{(k+1)!2^{\ell}}{(k-1-\ell)!(\ell+1)^{2}(2k-\ell)^{\ell+1}}.$
To evaluate the above expression, we aim to interchange the limit and the summation. To justify this step, we use the Dominated Convergence Theorem. Let $a_{k,\ell}$ denote the summand, defined as
$a_{k,\ell}=\frac{1}{k}\frac{(k+1)!2^{\ell}}{(k-1-\ell)!(\ell+1)^{2}(2k-\ell)^{\ell+1}}.$
Our goal is to show that
$\lim_{k\to\infty}\sum_{\ell=2}^{k-1}a_{k,\ell}=\sum_{\ell=2}^{\infty}\lim_{k\to\infty}a_{k,\ell}.$
To that end, we begin by analyzing the asymptotic behavior of $a_{k,\ell}$ as $k\to\infty$.
Using standard approximations, we observe that
$\frac{(k+1)!}{(k-1-\ell)!}\sim k^{\ell+2},\quad\text{and}\quad(2k-\ell)^{\ell+1}\sim(2k)^{\ell+1}.$
Therefore,
$\lim_{k\to\infty}a_{k,\ell}=\lim_{k\to\infty}\frac{2^{\ell}k^{\ell+2}}{(\ell+1)^{2}k(2k)^{\ell+1}}=\frac{1}{2(\ell+1)^{2}}.$ (20)
Next, we aim to find a summable sequence $M_{\ell}$ such that $|a_{k,\ell}|\leq M_{\ell}$ for all $k$.
From the approximation in20, we have
$a_{k,\ell}\sim\frac{1}{2(\ell+1)^{2}},$
which decays rapidly as $\ell\to\infty$. Moreover, for fixed $\ell$, the sequence $a_{k,\ell}$ is eventually decreasing in $k$.
Hence, there exists a constant $C$ (independent of $k$) such that
$|a_{k,\ell}|\leq\frac{C}{(\ell+1)^{2}}.$
Since $\sum_{\ell=2}^{\infty}\frac{1}{(\ell+1)^{2}}<\infty$, we conclude that $\{a_{k,\ell}\}$ is dominated by a summable sequence.
Thus, by the Dominated Convergence Theorem, we can interchange the limit and the summation:
$\lim_{k\to\infty}\sum_{\ell=2}^{k-1}a_{k,\ell}$ $=\sum_{\ell=2}^{\infty}\lim_{k\to\infty}a_{k,\ell}=\sum_{\ell=2}^{\infty}\frac{1}{2(\ell+1)^{2}}$
$=\frac{1}{2}\sum_{m=3}^{\infty}\frac{1}{m^{2}}=\frac{1}{2}(-1-\frac{1}{4}+\sum_{m=1}^{\infty}\frac{1}{m^{2}})$
$=\frac{1}{2}(\frac{\pi^{2}}{6}-\frac{5}{4})=\frac{\pi^{2}}{12}-\frac{5}{8}.$
Together with all the previous expressions, we obtain
$\liminf_{k\to\infty}\frac{T(k)}{k}\leq\frac{1}{2}+\frac{1}{8}+\frac{\pi^{2}}{12}-\frac{5}{8}=\frac{\pi^{2}}{12}\approx 0.822467.$
Figure 6 displays the normalized upper bound on $T(k)$ provided by Corollary 3.
20204040606080801001001201201401401601601801802002000.80.80.810.810.820.820.830.830.840.840.850.850.860.860.870.870.880.88Dimension $k$
Figure 6: Normalized (by $k$) upper bound on the random access expectation $T(k)$ for different values for $k$ from Corollary 3.